# Topological Center of the Double Dual of the Orlicz Figà-Talamanca Herz algebra

### v1

@@ Abstract.
For a locally compact group and a Young function let be the Orlicz analogue of the Figà-Talamanca Herz algebra Let denote the topological center of the double dual of when considered as a Banach algebra with the first Arens product. In this article, we present certain results concerning the existence of left and right identities in the Banach algebra and explore their relationship with the space Further, for amenable groups, we establish a necessary and sufficient condition for the equality to hold. Moreover, we derive several results related to the semi-simplicity of the Banach algebras and
@@
Primary 43A15, 43A25, 46J10; Secondary 43A07, 46E301. Introduction
@@
Let be a Banach algebra and let denote its double dual. In 1951, Arens [3] introduced the first and second Arens products on proving that is a Banach algebra when equipped with either of these products. In this article, we refer to the first Arens product as the Arens product, denoted by From the definition, it follows that the map is continuous on for any However, the map is not necessarily continuous. This observation leads to the definition of the topological center of . The set
is called the topological center of It is easy to verify that
@@
Let be the norm closure of the linear span of in Then, the space is also a Banach algebra when equipped with the multiplication induced by the Arens product on Over the past three to four decades, the study of the topological center and related problems in the algebras and has garnered significant interest. In 1987, Isik et al. [29] investigated the topological center problems for the group algebra and established that the topological center of the algebra is where is any locally compact group.
@@
In 1964, Eymard [15] introduced and studied the Fourier algebra and its dual for any locally compact group If the underlying group is abelian with as its dual group, then the Fourier algebra is isometrically isomorphic to In 1974, Granirer [23] defined the subspace of uniformly continuous functionals on denoted by When is abelian, the space is precisely the space of bounded uniformly continuous valued functions on the dual group
@@
Cecchini and Zappa [5] as well as Lau and Losert [35] investigated the topological center problems for the Banach algebra and the corresponding space Their results show that for a large class of locally compact groups including the motion group, the group and the Heisenberg group, the topological center of the algebra is and that of is where denotes the Fourier-Stieltjes algebra of In 1996, Lau and Ülger [37] further explored the topological center problems for the group algebra and the Fourier algebra yielding significant results.
@@
For and any locally compact group Figà-Talamanca [16] and Herz [27] introduced and studied the analogue of the Fourier algebra This algebra is referred to as the Figà-Talamanca Herz algebra and is denoted by The algebra of pseudomeasures, denoted is defined as the closure of in where denotes the representation of on given by left convolution. It is proved that the dual of is isometrically isomorphic to where and are dual Hölder exponents. For the algebra and its dual coincide with the Fourier algebra and the group von Neumann algebra respectively.
@@
The algebra has garnered significant attention from various researchers in recent decades, including Daws [10], Derighetti [12, 13], Forrest [17, 18], Granirer [24, 25, 26], Lau and Ülger [36] and Miao [38, 39, 40, 41], among others. The algebra of pseudomeasures has been explored in recent works by Daws and Spronk [11] and also by Gardella and Thiel [19, 20, 21, 22]. In 2009, Miao [41] investigated the topological center problems for the algebra and the corresponding space establishing a necessary and sufficient condition for the equality to hold.
@@
It is well known that the Orlicz spaces are an important generalization of the classical Lebesgue spaces and have attracted considerable attention. Recently, Lal and Kumar [30] as well as Aghababa and Akbarbaglu [1] independently introduced and studied the Orlicz analogue of the Figà-Talamanca Herz algebra by replacing the spaces with spaces, where is a Young function satisfying the condition. The version of the algebra is denoted by and is referred to as the Orlicz Figà-Talamanca Herz algebra. If is a complementary pair of Young functions, then the dual of is isometrically isomorphic to the space of pseudomeasures, denoted by [30, Theorem 3.5]. The algebra and its dual have received a great deal of attention in recent years; see [1, 2, 6, 7, 8, 30, 31, 32, 33].
@@
In this article, we focus on investigating the topological center problems for the algebra and the corresponding space The article is structured as follows: In Section 2, we recall the basic definitions, notations and some preliminary results. In Section 3, we present several results related to the existence of left and right identities in and explore their relationship with the space Theorem 3.4 characterizes the amenability of the group in terms of the existence of a right identity in assuming that the Young function satisfies the MA condition.
@@
In Section 4, Theorem 4.5 establishes a sufficient condition for any element in under which its restriction to the space belongs to This theorem plays a pivotal role in proving the main result of the article regarding the topological center of the double dual of In Section 5, we investigate the topological center problems for the Banach algebra and the corresponding space Theorem 5.2 provides a necessary and sufficient condition for the equality to hold, assuming that is amenable. This theorem serves as an Orlicz analogue of Miao’s result on algebras. Furthermore, an analogous result (Theorem 5.4) is established for the topological center of the Banach algebra
@@ 2. Preliminaries and Some Notations
This section begins by recalling the basic terminology associated with Orlicz spaces.
An even convex function is termed as a Young function if it satisfies the conditions and According to this definition, a Young function may take the value at certain points and hence be discontinuous at such points. However, in this context, we always assume that the Young function is both real-valued and continuous.
For each Young function a corresponding convex function is defined as follows:
This function is also a Young function and is referred to as the complementary function to Moreover, the pair (or ) is called a complementary pair of Young functions (or a complementary Young pair). As an example, for the function is a Young function, with its complementary function given by where and are dual Hölder exponents.
Let be a locally compact group with a left Haar measure A Young function is said to satisfy the condition (globally), denoted , if there exists a constant and such that the inequality
holds for all [42, Definition 1, Pg. 22]. By [42, Remark, Pg. 46], is said to be regular if the condition holds locally for when the measure is finite and globally when the measure is infinite. A pair of complementary Young functions is said to satisfy the condition if both and individually satisfy the condition. It is straightforward to verify that the family of Young functions for and satisfies the condition. Another example of such a function is However, its complementary function does not satisfy the condition.
A Young function is said to satisfy the Milnes-Akimoni condition (in short, MA condition) if for each there exists and such that
An example of a Young function satisfying the MA condition is for However, its complementary function for does not satisfy the MA condition. It is easy to verify that the family of Young functions with satisfies the MA condition. Furthermore, the family of complementary pairs of Young functions also satisfies the condition.
For a Young function the associated Orlicz space is denoted by and is defined as follows:
The Orlicz space is a Banach space when equipped with the Luxemburg (or Gauge) norm, defined by
Let be the complementary function to The Orlicz norm on is given by
In fact, these two norms are equivalent and for any the following holds:
If a Young function satisfies the condition, then the space of all continuous functions on with compact support is dense in Further, if the complementary function is continuous and satisfies if and only if then the dual space of is isometrically isomorphic to [42, Corollary 9, Pg. 111]. In particular, if the complementary Young pair satisfies the condition, then the Banach space is reflexive [42, Theorem 10, Pg. 112].
@@
Let be a Young function satisfying the condition. For any function , we define by for all The space is defined as the set of all continuous functions on that can be expressed in the form
@@
Note that if then For any the norm is given by
Equipped with this norm, along with pointwise addition and multiplication, the space becomes a commutative Banach algebra. The algebra is termed as the Orlicz Figà-Talamanca Herz algebra. Further, is a regular, tauberian, semi-simple Banach algebra, with its Gelfand spectrum homeomorphic to
Recall from [30, Lemma 4.1] that for any open subgroup of there exists an isometric isomorphism between and the subalgebra of consisting of functions that vanish outside We denote the inclusion map from to by and the restriction map from to by It follows from [30, Lemma 4.5] that the restriction map is a contractive homomorphism of Banach algebras. Further, it is easy to verify that is a closed ideal of using the aforementioned identification. Moreover, observe that the map is a bounded projection which is also a multiplier, i.e., it satisfies
@@
Each element defines a bounded linear operator where for all The space is a commutative Banach algebra, equipped with the operator norm and pointwise addition and multiplication. Furthermore, it is evident that
Let denote the set of all bounded complex Radon measures on For and we define the map by left convolution, i.e.,
If denotes the space of all bounded linear operators on equipped with the operator norm, it is straightforward to verify that Let be the closure of the set in with respect to the ultra-weak topology (the -topology). It is proved that for a locally compact group the dual of is isometrically isomorphic to [30].
Let be the norm closure of the set within The dual of is where is a commutative Banach algebra that contains [33]. For each we denote the operator by Note that represents the left regular representation of on
A linear functional is said to be a topologically invariant mean if it satisfies
@@
For further details on the Banach algebra and the related spaces, we refer the readers to the series of papers [1, 2, 6, 7, 8, 30, 31, 32, 33].
Let be a Banach algebra and let denote its double dual. It is known that can be endowed with two products which extend the multiplication of making a Banach algebra. These products were introduced by Arens [3] for normed algebras. In this article, we focus on the first Arens product, denoted by which is defined as follows:
Recall that, a Banach algebra is said to be weakly sequentially complete (WSC) if every weakly Cauchy sequence converges in the weak topology.
A locally compact group is said to be amenable if it admits a left-invariant mean, i.e., there exists a positive linear functional on with norm one such that
Examples of amenable groups include abelian groups, compact groups and solvable groups. However, the free group on two generators is not amenable.
Throughout this article, denotes a locally compact group with a fixed Haar measure and the pair of complementary Young functions satisfies the condition.
@@ 3. The Double Dual of and the Space
Here is the main theorem of this section, which characterizes the amenability of the group in terms of the existence of a right identity in the algebra This result generalizes Lau’s result [34, Proposition 3.2 (a)] on for any locally compact group Furthermore, the analogous result for hypergroups is given by Esmailvandi et al. [14, Proposition 4.1].
@@
Assume that is amenable. Then, by [32, Theorem 3.1], the algebra possesses a bounded approximate identity. Hence, by [4, Proposition III.28.7], there exists a right identity in the Banach algebra
Conversely, let be a right identity in Now, by Goldstine’s theorem ([9, Theorem A.3.29]), there exists a net in such that
@@
The result follows from [6, Theorem 3.8] and the fact that the algebra is semi-simple. ∎
@@
Further, by using the fact that the algebra is contained in the center of [8, Theorem 3.2], one can similarly prove that Thus, is an ideal in
Now, if possible, assume that and are two distinct elements of Then, by definition, which contradicts the fact that the Banach algebra is semi-simple. This implies that is singleton and hence, by [30, Corollary 7.6], is discrete. ∎
@@
Note that the Banach algebras and are semi-simple. This follows directly from the relation [33, Corollary 3.2] and the fact that the set is contained in the Gelfand spectrum of all of these algebras.
@@
Since is discrete, by [31, Corollary 4.13 (ii)], and thus, [33]. The desired result follows directly from Remark 3.8.
@@
Now, by repeating the arguments as in Theorem 3.6, it follows that is an ideal in and Since, by [30, Corollary 6.2], let be the restriction of to for some non-zero Observe that is also non-zero and belongs to Otherwise, if then
@@
This contradicts the fact that is non-zero and hence, Thus, by repeating the steps as in Theorem 3.6, it follows that is a singleton set and hence, by [30, Corollary 7.6], is discrete. ∎
@@ 4. Restriction Theorem
The aim of this section is to prove the restriction theorem (Theorem 4.5), which plays a pivotal role in establishing the main theorem of this article (Theorem 5.2), concerning the topological center of the double dual of The ideas behind the proofs are motivated by [41].
@@
If is an open subgroup of an amenable group then every right unit of in can be extended to a right unit of
@@
Recall from [28, Pg. 101] that for an algebra of functions the support of a linear functional as a subset of is characterized as follows: if and only if there exists a neighborhood of such that for all with
@@
Let and suppose that for every open compact subgroup of is in Then, for each there exists a compact subset of such that for all with compact support, and
@@
Let us suppose to the contrary that there exists and a function with compact support such that and where Now, by reiterating the construction outlined in [41, Lemma 3.1], we obtain a sequence of compactly supported functions in and a sequence of relatively compact symmetric neighborhoods of in such that, for each
@@
In the above computation, since the term equals zero [28, Pg. 101]. Now,
@@
Here is the main theorem of this section, which provides a sufficient condition for any under which the restriction of to belongs to The analogous result for the Fourier algebra where is a hypergroup and for Figà-Talamanca Herz algebra where is a locally compact group, are given by Esmailvandi et al. [14, Proposition 5.5] and Miao [41, Lemma 3.1], respectively. The following theorem is the Orlicz analogue of Miao’s result.
@@
Let If for every open compact subgroup of is in then the restriction of to is in
@@
for all with compact support and hence, by [31, Corollary 4.7], the restriction of to is The desired result is a direct consequence of the hypothesis that is in ∎

### v2

@@ Abstract.
Let a locally compact group and be a complementary pair of Young functions. Let be the Orlicz analogue of the classical Figà-Talamanca Herz algebra In this article, we establish a necessary and sufficient condition for the equality to hold, where denotes the topological center of the double dual of when equipped with the first Arens product. Furthermore, we prove several results concerning the semi-simplicity of the Banach algebras and
@@
Primary 43A15, 46J10, 46E30; Secondary 43A07, 43A251. Introduction
@@
Let be a Banach algebra and let denote its double dual. In 1951, Arens [4] introduced the first and second Arens products on proving that is a Banach algebra when equipped with either of these products. In this article, we refer to the first Arens product as the Arens product and denote it by From the definition, it follows that the map is continuous on for any In contrast, the map is not necessarily continuous. This asymmetry motivates the definition of the topological center of given by
It is straightforward to verify that
@@
Let be the norm closure of the linear span of in The dual space also becomes a Banach algebra when equipped with the multiplication induced by the Arens product on Over the past three to four decades, the study of the topological center and related aspects of the algebras and has been the subject of considerable research interest. In 1987, Isik et al. [30] addressed the topological center problems for the group algebra and established that the topological center of the algebra coincides with for any compact group
@@
Earlier, in 1964, Eymard [16] introduced and studied the Fourier algebra and its dual for any locally compact group If the underlying group is abelian with as its dual group, then the Fourier algebra is isometrically isomorphic to Later, in 1974, Granirer [24] defined the subspace of uniformly continuous functionals on denoted by When is abelian, this space coincides with the space of bounded uniformly continuous valued functions on the dual group
@@
Cecchini and Zappa [6] as well as Lau and Losert [36, 37], investigated the topological center problems for the Banach algebra and the corresponding space Their results show that, for a large class of locally compact groups including the abelian groups, discrete amenable groups, the motion group, the affine group and the Heisenberg group, the topological center of the algebra is and the topological center of is the Fourier-Stieltjes algebra In 1996, Lau and Ülger [39] extended this line of research further analyzing the topological center problems for both the group algebra and the Fourier algebra yielding significant results. In 2016, Losert [40] proved that if contains a free group then the topological center of is strictly larger than
@@
For and any locally compact group Figà-Talamanca [17] and Herz [28] introduced the -analogue of the Fourier algebra known as the Figà-Talamanca Herz algebra, denoted by The algebra of pseudomeasures, denoted is defined as the closure of in where is the representation of on given by left convolution. It is proved that the dual space of is isometrically isomorphic to where and are dual Hölder exponents. In the special case the algebra and its dual coincide with the Fourier algebra and the group von Neumann algebra respectively.
@@
The algebra has received significant attention in recent decades from numerous researchers, including Daws [11], Derighetti [13, 14], Forrest [18, 19], Granirer [24, 25, 26, 27], Lau and Ülger [38] and Miao [41, 42, 43, 44], among others. The corresponding algebra of pseudomeasures has also been the subject of extensive study, notably in the work of Daws and Spronk [12] as well as Gardella and Thiel [20, 21, 22, 23]. In 2009, Miao [44] investigated the topological center problems for the algebra and the corresponding space establishing a necessary and sufficient condition for the equality to hold.
@@
It is well known that the Orlicz spaces are an important generalization of the classical Lebesgue spaces and have attracted considerable attention. Recently, Lal and Kumar [31] and independently Aghababa and Akbarbaglu [1] introduced and studied the Orlicz -analogue of the Figà-Talamanca Herz algebra by replacing the spaces with spaces, where is a Young function satisfying the condition. The resulting algebra, denoted is known as the Orlicz Figà-Talamanca Herz algebra. If is a complementary pair of Young functions, then the dual space of is isometrically isomorphic to the space of pseudomeasures, denoted by [31, Theorem 3.5]. Both the algebra and its dual have received a great deal of interest in recent years; see [1, 2, 3, 7, 8, 9, 31, 32, 33, 34].
@@
In this article, we investigate the topological center problems for the algebra and the corresponding space extending certain known results to a broader functional-analytic framework. The structure of the article is as follows: Section 2 provides the necessary background, including basic definitions, notations and preliminary results. Section 3 focuses on the existence of left and right identities in and explore their relationship with the space In particular, Theorem 3.4 characterizes the amenability of the group in terms of the existence of a right identity in under the assumption that the Young function satisfies the MA condition.
@@
In Section 4, we prove Theorem 4.5, which provides a sufficient condition under which the restriction of an element in to lies in This result plays a central role in establishing the main theorem of the article regarding the topological center of the double dual of Section 5 is devoted to investigating the topological center problems for both the algebra and the corresponding space In Theorem 5.2, we establish a necessary and sufficient condition for the equality to hold, assuming that is amenable. This result can be seen as an Orlicz analogue of Miao’s theorem for algebras. Moreover, an analogous result (Theorem 5.4) is established for the topological center of the Banach algebra
@@ 2. Preliminaries and Some Notations
This section begins with a review of the fundamental terminology related to Orlicz spaces.
An even convex function is termed as a Young function if it satisfies the conditions and According to this definition, a Young function may attain the value at certain points and may therefore be discontinuous. However, in this context, we shall always assume that is real-valued and continuous.
For each Young function a corresponding convex function known as the complementary function to is defined by
The function thus defined is also a Young function. The pair (or equivalently ) is referred to as a complementary pair of Young functions, or a complementary Young pair. As an example, for the function is a Young function, and its complementary function is where and are dual Hölder exponents.
Let be a locally compact group equipped with a left Haar measure A Young function is said to satisfy the condition (globally), denoted , if there exists a constant and such that
for all [45, Definition 1, Pg. 22]. By [45, Remark, Pg. 46], is said to be regular if the condition holds locally for when the measure is finite and globally when the measure is infinite. A pair of complementary Young functions is said to satisfy the condition if both and individually satisfy the condition. It is straightforward to verify that the family of Young functions for and satisfies the condition. Another example is the function which satisfies the condition However, its complementary function does not satisfy the condition.
A Young function is said to satisfy the Milnes-Akimoni condition (abbreviated as the MA condition) if, for each there exists constants and such that
An example of a Young function satisfying the MA condition is for However, its complementary function for does not satisfy the MA condition. It is easy to verify that the family of Young functions defined by with satisfies the MA condition. Moreover, the corresponding family of complementary pairs of Young functions also satisfies the condition.
For a given Young function the associated Orlicz space is denoted by and is defined as
The space becomes a Banach space when equipped with the Luxemburg norm (also known as the Gauge norm), defined by
Let be the complementary function to Then, the Orlicz norm on is given by
In fact, these two norms are equivalent and for every the following holds:
If a Young function satisfies the condition, then the space of all continuous functions on with compact support is dense in Further, if the complementary function is continuous and satisfies if and only if then the dual space of is isometrically isomorphic to [45, Corollary 9, Pg. 111]. In particular, if the complementary Young pair satisfies the condition, then the Banach space is reflexive [45, Theorem 10, Pg. 112].
@@
Let be a Young function satisfying the condition. For any function , we define by for all The Orlicz Figà-Talamanca Herz algebra is defined as the set of all continuous functions that can be expressed as
@@
For any the norm is defined by
Equipped with this norm, along with pointwise addition and multiplication, the space becomes a commutative Banach algebra. Moreover, it is a regular, tauberian, semi-simple Banach algebra with its Gelfand spectrum homeomorphic to
Recall from [31, Lemma 4.1] that for any open subgroup of the Banach algebra is isometrically isomorphic to the subalgebra of consisting of functions that vanish outside We denote by the inclusion map from to and by the restriction map from to It follows from [31, Lemma 4.5] that the map is a contractive homomorphism of Banach algebras. Further, it is easy to verify that is a closed ideal of using the aforementioned identification. Moreover, the map is a bounded projection which is also a multiplier, i.e., it satisfies
@@
Each element induces a bounded linear operator given by for all Equipped with the operator norm and pointwise operations (addition and multiplication), the space is a commutative Banach algebra. Moreover, it is clear that
Let denote the set of all bounded complex Radon measures on For and we define the operator by left convolution:
If denote the Banach space of all bounded linear operators on equipped with the operator norm. It is straightforward to verify that for each Let be the closure of the set in with respect to the ultra-weak topology (the -topology). It is proved that the dual space of is isometrically isomorphic to [31].
Let denote the norm closure in of the set where is the measure associated with The dual of is denoted by which is a commutative Banach algebra containing [34]. For each the associated operator is denoted by The map thus defines the left regular representation of on
A linear functional is called a topologically invariant mean if it satisfies the following conditions:
@@
For additional details on the Banach algebra and the related spaces, we refer the readers to the series of papers [1, 3, 7, 8, 9, 31, 32, 33, 34].
Let be a Banach algebra and let denote its double dual. It is well known that can be equipped with two natural extensions of the multiplication on both of which turn into a Banach algebra. These extensions, introduced by Arens [4], are commonly referred to as the first and second Arens product. In this article, we focus on the first Arens product, denoted by which is defined as follows:
Recall that a Banach algebra is said to be weakly sequentially complete (WSC) if every weakly Cauchy sequence converges with respect to the weak topology.
A locally compact group is said to be amenable if there exists a left-invariant mean on i.e., a positive linear functional on of norm one satisfying
where denotes the left translation of Classical examples of amenable groups include abelian groups, compact groups and solvable groups. In contrast, the free group on two generators is a well known example of a non-amenable group.
Throughout this article, denotes a locally compact group with a fixed Haar measure and assume that the complementary pair of Young functions satisfies the condition.
@@ 3. The Double Dual of and the Space
Here is the main theorem of this section, which characterizes the amenability of the group in terms of the existence of a right identity in the algebra This result generalizes Lau’s result [35, Proposition 3.2 (a)] on for any locally compact group Furthermore, the analogous result for hypergroups is given by Esmailvandi et al. [15, Proposition 4.1].
@@
Assume that is amenable. Then, by [33, Theorem 3.1], the algebra possesses a bounded approximate identity. Hence, by [5, Proposition III.28.7], there exists a right identity in the Banach algebra
Conversely, let be a right identity in Now, by Goldstine’s theorem ([10, Theorem A.3.29]), there exists a net in such that
@@
The result follows from [7, Theorem 3.8] and the fact that the algebra is semi-simple. ∎
@@
Further, by using the fact that the algebra is contained in the center of [9, Theorem 3.2], one can similarly prove that Thus, is an ideal in
Now, if possible, assume that and are two distinct elements of Then, by definition, which contradicts the fact that the Banach algebra is semi-simple. This implies that is singleton and hence, by [31, Corollary 7.6], is discrete. ∎
@@
Note that the Banach algebras and are semi-simple. This follows directly from the relation [34, Corollary 3.2] and the fact that the set is contained in the Gelfand spectrum of all of these algebras.
@@
Since is discrete, by [32, Corollary 4.13 (ii)], and thus, [34]. The desired result follows directly from Remark 3.8.
@@
Now, by repeating the arguments as in Theorem 3.6, it follows that is an ideal in and Since, by [31, Corollary 6.2], let be the restriction of to for some non-zero Observe that is also non-zero and belongs to Otherwise, if then
@@
This contradicts the fact that is non-zero and hence, Thus, by repeating the steps as in Theorem 3.6, it follows that is a singleton set and hence, by [31, Corollary 7.6], is discrete. ∎
@@ 4. Restriction Theorem
The aim of this section is to prove the restriction theorem (Theorem 4.5), which plays a pivotal role in establishing the main result of this article (Theorem 5.2), concerning the topological center of the double dual of The ideas behind the proofs are motivated by [44].
@@
Let be an open subgroup of an amenable group Then every right unit of in admits an extension to a right unit of
@@
Recall from [29, Pg. 101] that for an algebra of functions the support of a linear functional as a subset of is characterized as follows: if and only if there exists a neighborhood of such that for all with
@@
Let and assume that for every open compact subgroup of is in Then, for each there exists a compact subset of such that for all with compact support, and
@@
Let us suppose to the contrary that there exists and a function with compact support such that and where Now, by reiterating the construction outlined in [44, Lemma 3.1], we obtain a sequence of compactly supported functions in and a sequence of relatively compact symmetric neighborhoods of in such that, for each
@@
In the above computation, since the term equals zero [29, Pg. 101]. Now,
@@
Here is the main theorem of this section, which provides a sufficient condition for any under which the restriction of to belongs to The analogous result for the Fourier algebra where is a hypergroup and for Figà-Talamanca Herz algebra where is a locally compact group, are given by Esmailvandi et al. [15, Proposition 5.5] and Miao [44, Lemma 3.1], respectively. The following theorem is the Orlicz analogue of Miao’s result.
@@
Let If for every open compact subgroup of is in then the restriction of to belongs to
@@
for all with compact support and hence, by [32, Corollary 4.7], the restriction of to is The desired result is a direct consequence of the hypothesis that is in ∎